# Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions

### v1

@@ IV-A Parameters of the General Construction
Let $k,k_{1}$ be integers with $k>k_{1}\geq 2g(F)-1$, and $P$ a rational place of $F$. Define
$\mathcal{L}^{*}_{k_{1}}(kP):=\{f|-k\leq v_{P}(f)<-k_{1}\}\cup\{0\}.$
Then $\mathcal{L}^{*}_{k_{1}}(kP)$ forms a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$.
Proof.
By the Riemann-Roch theorem, we obtain that $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions $\ell_{k}:=k-g(F)+1$ and $\ell_{k_{1}}:=k_{1}-g(F)+1$, respectively. Thus,
$\mathcal{L}^{*}_{k_{1}}(kP):=\mathcal{L}(kP)\setminus\mathcal{L}(k_{1}P)\cup\{0\}.$
For $g_{1},g_{2}\in\mathcal{L}^{*}_{k_{1}}(kP)$ and $a\in\mathbb{F}_{q}$, we observe that $g_{1}+g_{2}\in\mathcal{L}(kP)$ and $ag_{1}\in\mathcal{L}^{*}_{k_{1}}(kP)$. It remains to show $g_{1}+g_{2}\notin\mathcal{L}(k_{1}P)$.
Let $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$ be a basis for $\mathcal{L}(kP)$. Since $k_{1}\geq 2g-1$, the space $\mathcal{L}(k_{1}P)$ has a basis contained in $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$. Suppose that
$g_{1}+g_{2}=\sum\limits_{i=\ell_{k_{1}}+1}^{\ell_{k}}a_{i}\omega_{i}$
with $a_{i}\in\mathbb{F}_{q}$. It follows that $g_{1}+g_{2}\in\mathcal{L}(k_{1}P)$ implies $g_{1}+g_{2}=a_{1}\omega_{1}+\cdots+a_{\ell_{k_{1}}}\omega_{\ell_{k_{1}}}$. The linear independence of basis elements forces all coefficients to vanish, which means $g_{1}+g_{2}=0\in\mathcal{L}^{*}_{k_{1}}(kP)$. ∎
Theorem IV.1.
Suppose that $F^{\prime}/F$ is a Kummer extension of degree 2 with a totally ramified place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$. Let $s,k,k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
Proof.
The code $C_{sr,1}$ has length exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, which implies that the dimension of $C_{sr,1}$ simplifies to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$. Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$. Then the codeword of $L$ has sum-rank distance $d_{L}=\sum\limits_{i=1}^{s}d_{i}$. By Lemma IV.2, we have
$v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
for each $i$. Furthermore, Lemma IV.1 shows
$\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, we get $\det(\varepsilon_{P_{\alpha}}(L))\neq 0$. When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &f_{22}u,&f_{21}\end{bmatrix}$ (1)
where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then It follows from $\det(\varepsilon_{Q_{i}}(L))=0$ that $u-f_{*}^{2}$ belong to $F$ for a certain function $f_{*}$. Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish as $u\neq\omega^{2}$ for any $\omega\in F$. Therefore $\det(\varepsilon_{P_{\alpha}}(L))$ has at most $k$ zeros. This implies that
$\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore we have $d_{sr}\geq 2s-k$. This completes the whole proof. ∎
IV-B New Construction for Quadratic Extension
For a quadratic extension $F^{\prime}/F$, we provide an alternative construction by utilizing completely split places. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Consider a place $Q_{0}$ that splits completely in $F^{\prime}/F$ with $Q_{0,1}$ and $Q_{0,2}$ lying over $Q_{0}$. Let $G=nQ,G_{i}=nQ_{0,i}$ and $D=Q_{1}+\cdots+Q_{s}$. The evaluation map is constructed as follows:
$\text{ev}_{D,2}:\bigoplus_{i=1}^{2}\mathcal{L}(G_{i})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
The image of $\text{ev}_{D,2}$ forms a linear subspace of $\Pi$. Subsequently, the code $C_{sr,2}(D,G)$ is defined as the image. To analyze the parameters of $C_{sr,2}(D,G)$, we require the following lemma.
Lemma IV.4.
Let $Q\neq P$ be a rational divisor of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
@@ IV-B New Construction for Quadratic Extension
Furthermore, if there exists $f_{i}$ such that
@@
First, every term in the determinant expansion of $\varepsilon_{P}(L)$ must be contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
@@
Following a similar argument to Lemma IV.1, we obtain
@@
If $L(g)=0$, then
$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))<v_{Q_{1}}(\sigma_{1}(g)).$
@@
The subsequent result parallels Theorem IV.1, thus we omit the details of the proof.
@@
Suppose that $F^{\prime}/F$ is a function field extension of degree 2. Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1}|Q_{0}$ and $Q_{0,2}|Q_{0}$. Let $s,k,k_{1}$ be integers such that $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
@@ IV-C Examples from Elliptic Function Fields
An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ of genus 1. This serves as a canonical example since evaluation places can be selected from all $q+1$ places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$, limiting the code length to $q$. Nevertheless, this suffices to illustrate our framework.
In this section, we consider the case when $q$ is odd, as elliptic function fields over even characteristic are Artin-Schreier extensions. Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
@@
with $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ being a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
@@
For positive integers $k,k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
@@
where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}\in\mathbb{F}_{q}(x)/\mathbb{F}_{q}$. Through the Singleton bound combined with Theorems IV.1 and IV.2, we derive
@@
Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
@@
$\left(\begin{bmatrix}3,&1\\ -3,&3\end{bmatrix},\ldots,\begin{bmatrix}3,&1\\ 5,&3\end{bmatrix}\right)$
achieves the maximum sum-rank distance $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
@@
$\left(\begin{bmatrix}1,&0\\ 0,&1\end{bmatrix},\ldots,\begin{bmatrix}0,&0\\ 0,&2\end{bmatrix}\right)$
has sum-rank distance 8, attaining the lower bound.
@@
Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
@@ V Conclusion
In this paper, we conducted two families of $2\times 2$ sum-rank metric codes by exploiting quadratic Kummer extensions of algebraic function fields. By representing elements in the extension fields as matrices over the base field through explicit algebraic mappings, we rigorously determined the parameters of the proposed codes, including their minimum sum-rank distances and code rates. To demonstrate the practical applicability of our framework, we provided concrete instantiations over elliptic function fields.
To the best of our knowledge, these code constructions are novel in the context of algebraic-geometric (AG) codes for the sum-rank metric. Prior to this work, the only systematic study of AG codes in this metric was established by Berardini and Caruso [4], who proposed a general framework for such codes. However, our methodology differs fundamentally from theirs by explicitly exploiting the intrinsic algebraic structures of function field extensions—particularly through matrix representations derived from quadratic Kummer extensions—to control code parameters and optimize distance properties.
In this paper, we investigate Kummer extensions defined by the equation $y^{2}=u\quad\text{where }u\notin\{\omega^{2}\mid\omega\in F\setminus\mathbb{F}_{q}\}.$ This specific choice of non-square elements $u$ enables us to rigorously establish the non-degeneracy condition $\det(\varepsilon_{P}(L))\neq 0$ in the restricted case where $f_{1}=0$. The principal challenge in generalizing these constructions to $n\times n$ matrices stems from the need to preserve the non-vanishing determinant property $\det(\varepsilon_{P}(L))\neq 0\quad\text{for arbitrary }n\geq 2.$ We intend to address this dimensionality constraint through more flexible algebraic frameworks in subsequent work.
Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension $k-g(F)+1$, we observe the following approximate relationship:
@@
This raises an interesting problem: whether techniques developed for constructing MDS AG codes [11] can be adapted to the sum-rank metric setting to achieve MSRD AG codes. Following the approach for AG codes with narrow bounds (such as elliptic codes [11]), we propose this problem as an important objective for our subsequent research.

### v2

@@ IV-B Construction with Splitting Completely Rational Places
Let $P$ be a rational place of $F$. Suppose that $Q\neq P$ is a place of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ are the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
@@
Furthermore, if there exists an $f_{i}$ satisfying
@@
First, every term in the determinant expansion of $\varepsilon_{P}(L)$ is contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
@@
By an argument similar to that in Lemma IV.1, we obtain
@@
If $L(g)=0$, that is, $f_{1}\sigma_{1}(g)+f_{2}\sigma_{2}(g)=0$. Then by the strict triangle inequality and $v_{Q_{1}}(f_{1})\leq 0$, it follows that
$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))\leq v_{Q_{1}}(\sigma_{1}(g)).$
@@
The subsequent result is analogous to Theorem IV.1, so we omit proof details.
@@
Suppose that $F^{\prime}/F$ is a function field extension of degree 2. Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1},Q_{0,2}$ lying over $Q_{0}$. Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For distinct places $Q_{1},\ldots,Q_{s}$, the code
@@ IV-C Examples from Elliptic Function Fields
An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ and has genus 1. It serves as a canonical example and suffices to illustrate our framework.
In this section, we consider the case when $q$ is odd. Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
@@
where $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ is a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
@@
For positive integers $k$ and $k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
@@
where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}$. By applying the Singleton bound together with Theorems IV.1 and IV.2, we obtain
@@
Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ defined over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
@@
[…]
achieves the maximum sum-rank weight of $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
@@
[…]
has a sum-rank weight that attains the lower bound.
@@
Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}+P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
@@
IV-D The Gilbert-Varshamov-like Bound
In this section, we asymptotically compare our codes with the GV-like bound for sum-rank metric codes. Suppose that $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}/F_{i}$ is a Galois extension of degree 2 for all $i\in\mathbb{N}$. Denote by $\{g_{i}\}_{i\in\mathbb{N}}$ and $\{N(F_{i})\}_{i\in\mathbb{N}}$ the genus and the number of rational places of $F_{i}$, respectively. Following the construction from previous sections, with a sequence of integers $\{k_{i}\}_{i\in\mathbb{N}}$ such that $2g_{i}-1\leq k_{i}\geq 2N(F_{i})$, we obtain a sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ with parameters
  • code length: $s_{i}\leq 4N(F_{i})$,
  • dimension: $\ell_{k_{i}}=k_{i}-g_{i}+1$,
  • minimal sum-rank distance: $d_{sr,i}\geq 2s_{i}-k_{i}$.
Therefore, we have
$d_{sr,i}+\ell_{k_{i}}\geq 2s_{i}-g_{i}+1.$
By the Hasse-Weil lower bound over $\mathbb{F}_{q}$, if either $N(F_{i})$ or $g_{i}$ tends to infinity, then so does the other.
Ihara [20] observed that the Hasse-Weil bound is no longer optimal when the genus of a function field is sufficiently large with respect to $q$. Let $A(q)$ be the Ihara’s constant, defined as
[…]
The Drinfeld-Vlǎduţ bound [47] states that $A(q)\leq\sqrt{q}-1$. If $q$ is a square, then Drinfeld and Vlǎduţ proved that equality holds. Tsfasman, Vlǎduţ, and Zink [45] provided an example of curves that attain this bound employing modular curves. Garcia and Stichtenoth [14] presented a more explicit tower of function fields. Using their technology, we can present a chain of quadratic function field extensions that attain the bound for some even $q$.
Proposition IV.1.
Let $q=2^{2^{h}}$ with $h\geq 2$, and $F_{0}=\mathbb{F}_{q}(z_{0})$ be the rational function field. Suppose that $t=2^{h-1}$ and $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}=F_{i}(z_{i+1})$, where $z_{i}$ satisfies the equations:
  • for $j\geq 0$, $z_{jt+1}^{2}+z_{jt+1}=x_{j}^{2^{t}+1}$, where $x_{0}=z_{0}$, $x_{1}=z_{t}$ and $x_{j}=\frac{z_{jt}}{x_{j-1}}$ for $j\geq 2$,
  • for $j\geq 0$ and $2\leq i\leq t$, $z_{jt+i}^{2}+z_{jt+i}=z_{jt+i-1}$.
Then
$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\sqrt{q}-1.$
Proof.
It is sufficient to show that for any $j\geq 0$, the function field $F_{jt}$ has the form $F_{(j+1)t}=F_{jt}(z_{(j+1)t})$, where $z_{(j+1)t}$ satisfies the equation
$z_{(j+1)t}^{2^{t}}+z_{(j+1)t}=x_{j}^{2^{t}+1},$
with
$x_{0}=z_{0},\ x_{1}=z_{t},\ x_{j}=\frac{z_{jt}}{x_{j-1}}\text{ for }j\geq 2.$
Then, using [14, Corollary 3.2], we obtain
$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\lim\limits_{j\to\infty}\frac{N_{jt}}{g_{jt}}=\sqrt{q}-1.$
For $j=0$, observe that
$z_{t}^{2}+z_{t}=z_{t-1},$
$z_{t}^{4}+z_{t}^{2}=z_{t-1}^{2},$
$\cdots$
$z_{t}^{2^{t}}+z_{t}^{2^{t-1}}=z_{t-1}^{2^{t-1}}.$
Recall that $t=2^{h-1}$, which implies
$z_{t}^{2^{t}}+z_{t}$ $=(z_{t-1}^{2^{t-1}}+z_{t-1}^{2^{t-2}})+\cdots+(z_{t-1}^{2}+z_{t-1})$
$=z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-4}}+\cdots+z_{t-2}^{4}+z_{t-2}$
$=(z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-3}})+(z_{t-2}^{2^{t-3}}+z_{t-2}^{2^{t-4}})+\cdots+(z_{t-2}^{4}+z_{t-2}^{2})+(z_{t-2}^{2}+z_{t-2})$
$=z_{t-3}^{2^{t-3}}+z_{t-3}^{2^{t-4}}+\cdots+z_{t-3}^{2}+z_{t-3}$
$=z_{t-4}^{2^{t-4}}+z_{t-4}^{2^{t-6}}+\cdots+z_{t-4}^{4}+z_{t-4}$
$=z_{t-5}^{2^{t-5}}+z_{t-5}^{2^{t-6}}+\cdots+z_{t-5}^{2}+z_{t-5}$
$\qquad\vdots$
$=z_{1}^{2}+z_{1}$
$=x_{0}^{2^{t}+1}$
Then we have $F_{t}=F_{0}(z_{t})$ with
$z_{t}^{2^{t}}+z_{t}=x_{0}^{2^{t}+1}$
The proof can be completed with similar equations for any $j\geq 1$. ∎
The above proposition shows that there exist towers of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound for some square $q$. For any sum-rank metric code $C_{i}\in\mathbb{F}_{q}^{2\times 2}$ with $i\in\mathbb{N}$, its rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ are defined as
$\mathcal{R}_{i}:=\frac{\ell_{k_{i}}}{2s_{i}},\quad\delta_{sr,i}:=\frac{d_{sr,i}}{2s_{i}}.$
Theorem IV.3.
Suppose that $q$ is a square and there exists a tower $\{F_{i}\}_{i\in\mathbb{N}}$ of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound. For any real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
$\mathcal{R}_{sr}<1-\delta_{sr}-\frac{1}{2(\sqrt{q}-1)}.$ (1)
There exists a sum-rank metric code with rate at least $\mathcal{R}_{sr}$ and relative minimum distance at least $\delta_{sr}$.
Proof.
We only need to consider the sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ constructed from the tower $\{F_{i}\}_{i\in\mathbb{N}}$ via Theorem IV.1 or IV.2. For any $i\in\mathbb{N}$, the parameters of $C_{i}$ are
$s\leq 4N(F_{i}),\ell_{k_{i}},d_{sr,i}\geq 2s_{i}-k_{i}].$
Denote the rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ as above. Choose $k_{i}$ such that $\mathcal{R}_{i}\geq\mathcal{R}_{sr}$ for all $i$ and $\lim\limits_{i\to\infty}\mathcal{R}_{i}=\mathcal{R}_{sr}$. Then we have
$\mathcal{R}_{sr}=\lim\limits_{i\to\infty}\mathcal{R}_{i}\geq 1-\lim\limits_{i\to\infty}\delta_{sr,i}-\frac{1}{2(\sqrt{q}-1)}.$
Thus for sufficiently large $i$, the code $C_{i}$ satisfies the desired parameters. ∎
Note that the GV-like bound for $n=m=2$ has the form
$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{2}\log_{q}(1+\frac{1}{2\delta_{sr}})-\frac{\log_{q}(1+2\delta_{sr})}{4}$
$-\frac{\log_{q}(\gamma_{q})}{4}+o(1).$ (2)
For certain values of $q$ such that there exists a tower of quadratic function field extensions attaining the Drinfeld-Vlǎduţ bound, some families of our sum-rank metric codes surpass the GV-like bound. Using the function field tower given in Proposition IV.1, we plot the comparisons between equations (1) with (IV-D) for $q=2^{4}$, $q=2^{8}$ and $q=2^{16}$ in Figs. 2, 3, and 4, respectively.
Figure 2: $q=2^{4}$
Figure 3: $q=2^{8}$
Figure 4: $q=2^{16}$
@@ V Conclusion
In this paper, we presented two constructions of $2\times 2$ sum-rank metric codes via quadratic Galois extensions of algebraic function fields. By expressing the functions of extension fields to matrices over the base fields, we determined the parameters of these codes. Furthermore, we provided illustrative examples based on elliptic function fields.
To the best of our knowledge, these constructions are new. The only existing result on AG codes in the sum-rank metric was given by E. Berardini and X. Caruso [4]. In contrast to their work, our approach relies more heavily on the structures of function field extensions. In this paper, we established that $\det(\varepsilon_{P}(L))\neq 0$ in the special case where $f_{1}=0$. The primary challenge in extending these constructions to $n\times n$ matrices lies in guaranteeing $\det(\varepsilon_{P}(L))\neq 0$. Through SageMath computations, we verify that $L(g)$ has trivial kernel for certain special function fields and places when $n\geq 3$. Consequently, in our future work, we will investigate more general constructions.
Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension kg(F)+1k-g(F)+1, we have
@@
This raises an interesting problem: adapting the techniques used in constructing MDS AG codes to the sum-rank metric to obtain MSRD AG codes. Building on the approach for AG codes with narrow bounds (such as elliptic codes [11]), we identify this problem as an important objective for our subsequent research.